OpenStudy (anonymous):

help me with the steps. Find the center, transverse axis, vertices, foci, and asymptotes. x^2-9y^2-4x-72y-144=0

OpenStudy (tkhunny):

You must complete the square in x and y, separately. (x^2 - 4x + ____) - 9(y^2 + 8y + ____) = 144 + ____ + 9(____)

OpenStudy (anonymous):

so they will look like (x^2 - 4x + 4) -9(y^2 + 8y + 16 ) =144 + 4 + 144

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

That was very good. Did you need me to do all that setup or could you have done that yourself?

OpenStudy (anonymous):

didnt know how to put it in the fraction form and also forgot that need to add to the sum too. so how do I put in the form to find the center?

OpenStudy (tkhunny):

btw That should have been NEGATIVE 9 on the right hand side. Sorry about that. 144 + 4 - 144 = 4

OpenStudy (tkhunny):

Rewrite the two perfect square trinomials into their square binomial forms and divide by 4. You'll almost be there.

OpenStudy (anonymous):

so it will look like (x-2)^2 / 4 - 9(y+4)^2 / 4 =1

OpenStudy (anonymous):

and the center would be (2, -4)

OpenStudy (anonymous):

but not sure if it would be transverse on the x or y

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Write it like this. \(\dfrac{(x-2)^{2}}{4} - \dfrac{(y+4)^{2}}{4/9} = 1\)

OpenStudy (anonymous):

so it will make it transverse on the x

OpenStudy (tkhunny):

1) Wipe out the entire y-term. Is there a solution for x? Yes. This gives the direction and location of the vertices. Center: (2,-4) Vertex: (2 - 2,-4) and (2 + 2,-4) 2) On the other hand, wipe out the entire x-term. Is there a solution for y? No. This gives the direction and location of the "other" vertices. Center: (2,-4) Vertex: (2,-4 + 2/3) and (2,-4 - 2/3)

OpenStudy (anonymous):

so its transverse on the x since there is a solution for x and not for y

OpenStudy (tkhunny):

Yes.

OpenStudy (anonymous):

so the vertex you had how did you get them.

OpenStudy (tkhunny):

In that final form, the length of the transverse axis is pretty obvious. 4 = 2^2. It's that base 2 that's important. That's half the length of the transverse axis. That's why I inserted "+2" and "-2" to identify the vertices.

OpenStudy (anonymous):

oh okay that just confused me. to find the foci I would just need to do c^2 = a^2 +b^2 right so it will be c^2 = 4 + 4/9 and it will make c^2 = 24/5 and c will equal to sqroot of 24/5

OpenStudy (tkhunny):

?? Better try that addition again. 4 + 4/9

OpenStudy (anonymous):

oops i mean 40/9 for c^2

OpenStudy (tkhunny):

:-)

OpenStudy (anonymous):

so the foic will be (-sqroot 40/9, -4) and sqroot 40/9, -4)

OpenStudy (tkhunny):

1) I would prefer (2/3)sqrt(10) 2) No, that is (h+c,k) (h-c,k) You're 'h' is missing.

OpenStudy (anonymous):

so it will be (2-(2/3)sqroot 10, -4) and 2+(2/3)sqroot 10, -4)

OpenStudy (tkhunny):

I'd go with that! Good work. Asymptotes?

OpenStudy (anonymous):

to find asymtptes is it y=a/b x and y= -a/b x so it will be +- 3

OpenStudy (tkhunny):

Not quite. h and k are missing again. Those are good only for a center at the Origin. (y-k) = (b/a)(x-h) Ringing bells? One would also need to remember which way 'a' goes. You can just do it yourself, without the silly formula. One point on the line is (h,k) and the slope is, indeed, b/a. Just write it down from the point-slope form of a line.

OpenStudy (anonymous):

so its nto a/b but b/a and what is x?

OpenStudy (anonymous):

not*

OpenStudy (tkhunny):

If 'a' is horizontal, it's b/a. If 'a' is vertical, it's a/b. 'x' is part of the equation. Are you trying to replace it?

OpenStudy (anonymous):

oh no! nvm just notice that haha. and okay i understand now! thank you so MUCH for taking your time and helping me!